# The contact process on dynamical random trees with degree dependence

### v1

@@ 6 Proofs in the general graph setting
$\mathbb{P}(|\mathbf{C}_{t}|\geq L)\leq\frac{e^{\theta t}f(X)}{L}.$ (6.8)
@@
$\mathbb{P}(|\mathbf{C}_{t}|=\infty)=\lim_{L\to\infty}\mathbb{P}(|\mathbf{C}_{t}|\geq L)\leq\lim_{k\to\infty}\frac{e^{\vartheta t}f(X)}{L}=0.$
@@
For the last statements, we assume that $\vartheta<0$. From (6.7) and Markov’s property it follows that for each $n\in\mathbb{N}$ the process $M^{n}$ is a supermaringale such that $M_{t}^{n}\to 0$ as $t\to\infty$ almost surely. In addition, by (6.8) and the continuity of probability measures
@@
Now Theorem 3.1 follows as a corollary. We just need to pick a $\lambda>0$ small enough such that (6.6) is satisfied.
@@
Assume that $p$ and $v$ satisfy (2.3) with $\eta,\alpha\geq 0$. Since $p$ is decreasing in both parameters and $d_{x}\geq 1$ for any $x\in V$, we have
@@
With this in hands, we can check conditions (3.2) and (3.3) of Theorem 3.1. We first prove the statement under condition $(i)$. Therefore, we assume $\alpha\geq 1$. Then we choose $W(d_{x})=d_{x}$, then it follows that for all $x\in V$
@@
$\sum_{\{x,y\}\in E}W(d_{y})p_{\alpha,\sigma}(d_{x},d_{y})\leq\kappa\sum_{\{x,y\}\in E}d_{y}^{\beta}\big{(}(d_{x}\wedge d_{y})^{\sigma}(d_{x}\vee d_{y})\big{)}^{-\alpha}\leq\kappa\sum_{\{x,y\}\in E}d_{x}^{-\alpha\sigma}.$
It follows that for any $x\in V$
$\sum_{\{x,y\}\in E}W(d_{y})p_{\alpha,\sigma}(d_{x},d_{y})\leq\kappa d_{x}^{1-\alpha\sigma}\leq\kappa W(d_{x}),$
@@
where in the last inequality we used that $\alpha+2\eta\geq 1$. Therfore, again the conditions of Theorem 3.1 are satisfied, which yields the claim. ∎
@@
In this proof we denote by $\lambda_{2}^{p}=\lambda_{2}^{p}(\kappa,\sigma)$ the critical infection rate for local survival for the penalised contact process on the BGW tree with transitions as in (3.4) and with penalisation chosen to be $p_{\alpha,\sigma}$ as in (2.5) with $\kappa>0$, $\sigma\in[0,1]$ and $\alpha$ such that (3.9) holds. Note that for $\kappa=1$, $p_{\alpha,1}=(d_{x}d_{y})^{-\alpha}$ and by [1, Theorem 2.1] it follows that $\lambda_{2}^{p}(1,1)=0$. Now if $\kappa<1$, then $p_{\alpha,1}=\kappa(d_{x}d_{y})^{-\alpha}$, and thus it follows directly that $\lambda_{2}^{p}(\kappa,1)=\frac{1}{\kappa}\lambda_{2}^{p}(1,1)=0$. Furthermore, we see that $p_{\alpha,1}\leq p_{\alpha,\sigma}$ for all $\sigma\in[0,1]$, and this is also true if we replace $\kappa$ in $p_{\alpha,\sigma}$ with $\kappa^{\prime}$ such that $\kappa^{\prime}\geq\kappa$. Therefore, by monotonicity of the penalised contact process in the infection rates it follows that $0=\lambda_{2}^{p}(\kappa,1)\geq\lambda_{2}^{p}(\kappa^{\prime},\sigma)\geq 0$ This implies $\lambda_{2}^{p}(\kappa,\sigma)=0$ for all $\kappa>0$ and $\sigma\in[0,1]$. Finally, the claim that the critical infection rates of the CPDG with connection probability $p_{\alpha,\sigma}$ equal 00, i.e. $\lambda_{1}=\lambda_{2}=0$ follows by Proposition 3.3. ∎
@@ 7 Proofs in the tree setting
Due to Condition (2.3), for given $\alpha>0$ and $\eta\in\mathbb{R}$, we find constants $\kappa_{1},\kappa_{2},\nu_{1},\nu_{2}>0$ which only depend on $L$ such that
@@ 7.1 Stars and stable stars
As described in Subsection 3.2, a star can be thought of as reservoir for the infection, where it can survive for an exceptionally long time without input from outside. Furthermore, during this period we want to push the infection to the next star. Thus, for technical reasons it is more convenient if the path towards the next star and the reservoir are defined on disjoint parts of the graph, except for $x$. Thus, we need to be able to exclude a given vertex from the set of good neighbours. Therefore, we define for every $z\in\mathcal{N}$ the set
$\mathcal{G}_{k}(z)=\mathcal{G}_{x,k}(z):=\mathcal{G}_{x,k}\backslash\{z\},\quad k\geq 0.$
In the next result we show that with high probability we have enough good neighbours around the root $\rho$ for a sufficiently long time. This lemma can be proved analogously to [11, Proposition 3]. We will nevertheless give the whole proof for two reasons. First we adapted the statement slightly and secondly since we consider a different model compared to [11] there are minor differences in the proof, and thus we would like to highlight this once.
@@
There exists $c_{L}>0$ independent of $N,p(\cdot,\cdot)$ and $\lambda$ such that for $N$ large enough
$\mathbb{P}_{N}\big{(}|\mathcal{G}_{\rho,k}|-1\geq c_{L}Np(N,L)\text{ for all }k\leq e^{c_{L}Np(N,L)}\big{)}\geq 1-e^{-c_{L}Np(N,L)}.$
@@
As $\rho$ is fixed, we write $\mathcal{G}_{k}$ for $\mathcal{G}_{\rho,k}$ throughout. Note that for $y\in\mathcal{N}$ the random variable $|(\Delta^{\text{rec}}_{y}\cup\Delta^{\text{up}}_{\{\rho,y\}})\cap J_{k}|$ has a Poisson distribution with mean bounded by $44T(1+\nu_{2}N^{\eta})=$ 4, where we used that we conditioned on $D_{\rho}=N$, (7.1) and that by definition $1\leq D_{y}\leq L$. Now, by definition of the set $\mathcal{G}_{k}$ and $p(\cdot,\cdot)$ being a decreasing function, we have for every $y\in\mathcal{V}$ with $\{\rho,y\}\in\mathcal{E}$ that
@@
where we also used that we condition on $D_{\rho}=N$. Thus, we get that
$\mathbb{P}_{N}\big{(}y\in\mathcal{G}_{k}\big{)}\geq\phi_{L}e^{-4}p(N,L),$
where $\phi_{L}:=\mathbb{P}(\zeta_{y}\leq L-1)$, and we recall that $\zeta_{y}$ is the number of offspring of vertex $y$ and by (3.5) we have $D_{y}=\zeta_{y}+1$. Since the number of offspring are sampled independently for every parent in a BGW tree, this implies that $|\mathcal{G}_{k}|$ stochastically dominates a binomial random variable $B\sim\text{Bin}(N,e^{-4}\phi_{L}p(N,L)))$. By the Chernoff bound for a binomial distribution (see e.g. Theorem 2.21 in [7]), we get for $10<c<$ 1 that
@@
$\mathbb{P}_{N}\big{(}|\mathcal{G}_{k}|\geq ce^{-4}\phi_{L}$ $Np(N,L)\ \text{ for all }\ k\leq e^{ce^{-4}\phi_{L}Np(N,L)}\big{)}$ (7.3)
$\geq 1-\sum_{k=0}^{S}\mathbb{P}_{N}\big{(}|\mathcal{G}_{k}|<ce^{-4}\phi_{L}Np(N,L)\big{)}.$
@@
Note that $(1-c)^{2}-2c>2c$ if $c<3-2\sqrt{2}$, and thus if we choose $c\in(0,3-2\sqrt{2})$ and set $c_{L}:=ce^{-4}\phi_{L}/2$ we have
$\sum_{k=0}^{S}\mathbb{P}_{N}(|\mathcal{G}_{k}|<2c_{L}Np(N,L))\leq e^{-2c_{L}Np(N,L)}.$
Now for $N$ large enough it holds that $c_{L}Np(N,L)\geq 1$, and thus we have by plugging back into (7.3)
$\mathbb{P}_{N}\big{(}|\mathcal{G}_{k}|-1\geq c_{L}Np(N,L)\text{ for all }k\leq e^{c_{L}Np(N,L)}\big{)}$
$\geq\mathbb{P}_{N}\big{(}|\mathcal{G}_{k}|\geq 2c_{L}Np(N,L)\text{ for all }k\leq e^{c_{L}Np(N,L)}\big{)}$
$\geq 1-e^{-2c_{L}Np(N,L)}\geq 1-e^{-c_{L}Np(N,L)}$
which yields the claim. ∎
We say that $x$ is a stable star if the following event holds
$\overline{\mathcal{S}}=\overline{\mathcal{S}}_{x,N}:=\{|\mathcal{G}_{x,k}|-1>c_{L}Np(N,L)\ \text{for all}\ k\leq e^{c_{L}Np(N,L)}\}.$ (7.4)
Since we only consider stable stars in the subsequent subsection we mostly refer to $x$ just as a star. Note that Lemma 7.1 tells us that for $N$ large enough
$\mathbb{P}_{N}\big{(}\overline{\mathcal{S}}_{\rho,N})\geq 1-e^{-c_{L}Np(N,L)}.$
Furthermore, in most cases we need to exclude a specific neighbour $z\in\mathcal{N}_{x}$, and therefore we define the event
$\mathcal{S}(z)$ $=\mathcal{S}_{x,N}(z):=\{|\mathcal{G}_{x,k}(z)|>c_{L}Np(N,L)\ \text{for all}\ k\leq e^{c_{L}Np(N,L)}\}.$ (7.5)
The reason for excluding a specific neighbour $z\in\mathcal{N}_{x}$ is to ensure that the previously mentioned reservoir around $x$ and the path towards the next star only share the centre $x$ and otherwise are defined on disjoint parts of the graph. Note that $\overline{\mathcal{S}}$ acts as a lower bound since $\overline{\mathcal{S}}\subset\mathcal{S}(z)$ for all $z\in\mathcal{N}_{x}$.
@@ 7.2 Survival on a star
$\mathcal{G}_{k}^{\prime}(z)$ :=$:=\{y\in\mathcal{G}_{k}(z):\ y\in\mathbf{C}^{*}_{kT}\},$ (7.6)
where some $z\in\mathcal{N}_{\rho}$ is excluded. In words this is the set of good neighbours in the interval $J_{k}$, which are infected at time $kT$, and thus are infected at any time during $[kT,(k+2)T]$ since good neighbours do not recover during $J_{k}$. Furthermore, we denote for any $0<s<t$ and $z\in\mathcal{N}_{\rho}$ by $\mathcal{F}_{s,t}=\mathcal{F}_{s,t}(z)$ the $\sigma$-algebra generated by the graphical construction up to time $s$, and the processes $\Delta^{\text{rec}}_{y}$ and $\Delta^{\text{up}}_{\{\rho,y\}}$ for all $y\in\mathcal{N}_{\rho}\backslash\{z\}$ up to time $t$.
Note that in all results in the subsequent subsection we are only concerned with the number of good infected neighbours rather than which neighbours specifically are of this type or which neighbour exactly was excluded. Therefore, for the sake of readability, we omit the argument $z\in\mathcal{N}_{\rho}$, and write $\mathcal{G}_{k}$, $\mathcal{G}_{k}^{\prime}$, $\mathcal{S}_{\rho,N}$, and $\mathcal{F}_{s,t}$, as long as it is clear from the context.
In the next step we need to know how long the centre of a star is typically infected if we already have “enough” infected good neighbours. The following lemma states that the centre of a star with enough infected good neighbours will be infected for at least half of the time within a time interval $[kT,(k+1)T]$ with high probability.
@@
Let $N\in\mathbb{N}$. There exists a constant $C>0$ independent of $\lambda$, $T$ such that for any $M\in\mathbb{N}$ with $W:=\lambda TM>C$ and any $k\in\mathbb{N}$ we have
$\mathbb{P}_{\mathcal{T},N}\left(\int_{kT}^{(k+1)T}\mathds{1}_{\{\rho\notin\mathbf{C}_{s}^{*}\}}\mathrm{d}s>\frac{T}{2}\ \Big{|}\Big{.}\ \mathcal{F}_{kT,(k+2)T}\right)\mathds{1}_{\{|\mathcal{G}_{k}^{\prime}|\geq M\}}\leq Ce^{-W/4}.$
Proof.
The claim can be proved analogously as [11, Lemma 2]. ∎
$\mathcal{W}_{k}(z):=\left\{|\mathcal{G}^{\prime}_{k}(z)|\geq\delta\lambda Np(N,L)T\quad\text{and}\quad\int_{kT}^{(k+1)T}\mathds{1}_{\{\rho\notin\mathbf{C}_{s}^{*}\}}\mathrm{d}s\leq\frac{T}{2}\right\},$
where $z\in\mathcal{N}_{\rho}$. Here again we drop the dependence on $z$ and only write $\mathcal{W}_{k}$, as long as it is clear by context. Note again that the subsequent result is independent of the specific choice of $z$.
Lemma 7.3.
(Local survival) Let $\lambda>0$ and $N\in\mathbb{N}$ be chosen such that $\tfrac{3}{2}\lambda T_{N}<1$ and set $\bar{k}=\lfloor e^{\delta\lambda^{2}T^{2}Np(N,L)/4}\rfloor$. Then there exists a large universal constant $C$ such that if $\lambda^{2}T^{2}N>C$ for any $k\leq e^{c_{L}Np(N,L)/2}$, we have
@@
The result follows analogously to [11, Proposition 4]. ∎
With this result we have control over the event that an infection persists for an exponential time if we already have a sufficiently large pool of infected good neighbours which can sustains the centre to be infected most of the time. It remains to estimate the probability to reach this amount of good infected neighbours in case that only the centre is initially infected.
@@
Let $\lambda>0$ and $N\in\mathbb{N}$ be chosen such that $12\lambda T_{N}<$ 1. Furthermore, let $s\leq e^{c_{L}Np(N,L)/2}$ and $k=\lfloor s/T\rfloor$. Then it holds that
@@
By the choice of $k$ we see that $kT\leq s\leq(k+1)T$, and thus the probability that $\rho$ does not recover in $[s,(k+2)T]$ is bounded from below by $e^{-2T}$. On the other hand, the probability that an infection event happens between a good neighbour $y\in\mathcal{G}_{k+2}$ and $\rho$ in the time interval $[s,(k+2)T]$ is bound from below by $1-e^{-\lambda T}\geq\frac{\lambda T}{2}$. Note that the last inequality holds since $\lambda T<1$. Now given that $\rho$ does not recover in $[s,(k+2)T]$ we concluded that on the event $\mathcal{S}_{\rho,N}\cap\{\rho\in\mathbf{C}^{*}_{s}\}$ the random variable $|\mathcal{G}^{\prime}_{k+2}|$ is stochastically dominated by a binomial random variable $B$ with $\lceil 8\delta Np(N,L)\rceil$ trials and success probability $\tfrac{\lambda T}{2}$. Recall that $\delta=c_{L}/8$, where $c_{L}$ is the constant from Lemma 7.1. Therefore, we get that
@@
Next we can again use a Chernoff bound for the binomial distribution (see e.g. [7, Theorem 2.21]) to obtain
$\mathbb{P}_{\mathcal{T},N}\Big{(}B>\lceil 8\delta Np(N,L)\rceil\frac{\lambda T}{2}+\delta\lambda TNp(N,L)-\lceil 8\delta Np(N,L)\rceil\frac{\lambda T}{2}\Big{)}$
$\hskip 28.45274pt\geq 1-\exp\Big{(}-\frac{\lambda T(\lceil 8\delta Np(N,L)\rceil-2\delta Np(N,L))^{2}}{4\lceil 8\delta Np(N,L)\rceil}\Big{)}$
$\hskip 28.45274pt\geq 1-\exp\Big{(}-\frac{\lambda T(\lceil 8\delta Np(N,L)\rceil-4\delta Np(N,L))}{4}\Big{)}\geq 1-e^{-\delta\lambda TNp(N,L)},$
where in the last two inequalities we have used $(a-b)^{2}\geq a^{2}-2ab$ and $\lceil x\rceil\geq x$ for $x\geq 0$, respectively. Thus, it follows that
$\mathbb{P}_{\mathcal{T},N}\big{(}|\mathcal{G}^{\prime}_{k+2}|>\delta\lambda TNp(N,L)\mid\mathcal{F}_{s,(k+4)T}\big{)}\geq e^{-2T}(1-e^{-\delta\lambda TNp(N,L)}),$
as claimed. ∎

### v2

@@ 6 Proofs in the general graph setting
$\mathbb{P}(|\mathbf{C}_{t}|\geq L)\leq\frac{e^{\vartheta t}f(X)}{L}.$ (6.8)
@@
$\mathbb{P}(|\mathbf{C}_{t}|=\infty)=\lim_{L\to\infty}\mathbb{P}(|\mathbf{C}_{t}|\geq L)\leq\lim_{L\to\infty}\frac{e^{\vartheta t}f(X)}{L}=0.$
@@
For the last statements, we assume that $\vartheta<0$. From (6.7) and Markov’s property it follows that for each $n\in\mathbb{N}$ the process $M^{n}$ is a supermartingale such that $M_{t}^{n}\to 0$ as $t\to\infty$ almost surely. In addition, by (6.8) and the continuity of probability measures
@@
Proof of Theorem 3.1.
The claim of the theorem follows directly from Proposition 6.3. We just need to pick a $\lambda>0$ small enough so that (6.6) is satisfied. ∎
@@
Assume that $p$ and $v$ satisfy (2.3) and (2.4) with $\eta,\alpha\geq 0$. Since $p$ is decreasing in both parameters and $d_{x}\geq 1$ for any $x\in V$, we have
@@
With this in hand, we can check conditions (3.2) and (3.3) of Theorem 3.1. We first prove the statement under condition $(i)$. Therefore, we assume $\alpha\geq 1$. Then we choose $W(d_{x})=d_{x}$, then it follows that for all $x\in V$
@@
$\sum_{\{x,y\}\in E}W(d_{y})p_{\alpha,\sigma}(d_{x},d_{y})$ $\leq$ $\kappa\sum_{\{x,y\}\in E}d_{y}^{\beta}\big((d_{x}\wedge d_{y})^{\sigma}(d_{x}\vee d_{y})\big)^{-\alpha}$
$\leq$ $\kappa\sum_{\{x,y\}\in E}d_{x}^{-\alpha\sigma}\leq\kappa d_{x}^{1-\alpha\sigma}\leq\kappa W(d_{x}),$
@@
where in the last inequality we used that $\alpha+2\eta\geq 1$. Therefore, again the conditions of Theorem 3.1 are satisfied, which yields the claim. ∎
@@
In this proof we denote by $\lambda_{2}^{p}=\lambda_{2}^{p}(\kappa,\sigma)$ the critical infection rate for strong survival for the penalised contact process on the BGW tree with transitions as in (3.4) and with penalisation chosen to be $p_{\alpha,\sigma}$ as in (2.6) with $\kappa>0$, $\sigma\in[0,1]$ and $\alpha$ such that (3.9) holds. Note that for $\kappa=1$, $p_{\alpha,1}=(d_{x}d_{y})^{-\alpha}$ and by [1, Theorem 2.1] it follows that $\lambda_{2}^{p}(1,1)=0$. Now if $\kappa<1$, then $p_{\alpha,1}=\kappa(d_{x}d_{y})^{-\alpha}$, and thus it follows directly that $\lambda_{2}^{p}(\kappa,1)=\frac{1}{\kappa}\lambda_{2}^{p}(1,1)=0$. Furthermore, we see that $p_{\alpha,1}\leq p_{\alpha,\sigma}$ for all $\sigma\in[0,1]$, and this is also true if we replace $\kappa$ in $p_{\alpha,\sigma}$ with $\kappa^{\prime}$ such that $\kappa^{\prime}\geq\kappa$. Therefore, by monotonicity of the penalised contact process in the infection rates it follows that $0=\lambda_{2}^{p}(\kappa,1)\geq\lambda_{2}^{p}(\kappa^{\prime},\sigma)\geq 0.$ This implies $\lambda_{2}^{p}(\kappa,\sigma)=0$ for all $\kappa>0$ and $\sigma\in[0,1]$. Finally, the claim that the critical infection rates of the CPDG with connection probability $p_{\alpha,\sigma}$ equal 0, i.e. $\lambda_{1}=\lambda_{2}=0$ follows by Proposition 3.3. ∎
@@ 7 Proofs in the tree setting
Due to Conditions (2.3) and (2.4), for given $\alpha>0$ and $\eta\in\mathbb{R}$, we find constants $\kappa_{1},\kappa_{2},\nu_{1},\nu_{2}>0$ which only depend on $L$ such that
@@ 7.1 Stars and stable stars
As described in Subsection 3.2, a star can be thought of as reservoir for the infection, where it can survive for an exceptionally long time without input from outside. Furthermore, during this period we want to push the infection to the next star. In the next result we show that with high probability we have enough good neighbours around the root $\rho$ for a sufficiently long time. This lemma can be proved analogously to [11, Proposition 3.7]. We will nevertheless give the whole proof for two reasons. First we adapted the statement slightly and secondly since we consider a different model compared to [11] there are minor differences in the proof, and thus we would like to highlight this once. Before stating the result, we say that $x$ is a stable star if the following event holds
$\mathcal{S}=\mathcal{S}_{x,N}:=\{|\mathcal{G}_{x,k}|>c_{L}Np(N,L)\ \text{for all}\ k\leq e^{c_{L}Np(N,L)}\}.$ (7.2)
Since we only consider stable stars in the subsequent subsection we mostly refer to $x$ just as a star.
@@
There exists $c_{L}\in(0,1)$ independent of $N,p(\cdot,\cdot)$ and $\lambda$ such that for $N$ large enough
$\mathbb{P}_{N}\big(\mathcal{S}_{x,N}\big)\geq 1-e^{-c_{L}Np(N,L)}.$
@@
As $\rho$ is fixed, we write $\mathcal{G}_{k}$ for $\mathcal{G}_{\rho,k}$ throughout. Note that for $y\in\mathcal{N}$ the random variable $|(\Delta^{\text{rec}}_{y}\cup\Delta^{\text{up}}_{\{\rho,y\}})\cap J_{k}|$ has a Poisson distribution with mean bounded by $4T(1+\nu_{2}N^{\eta})=4$, where we used that we conditioned on $D_{\rho}=N$, (7.1) and that by definition $1\leq D_{y}\leq L$. Now, by definition of the set $\mathcal{G}_{k}$ and $p(\cdot,\cdot)$ being a decreasing function, we have for every $y\in\mathcal{V}$ with $\{\rho,y\}\in\mathcal{E}$ that
@@
where we also used that we condition on $D_{\rho}=N$. Thus, we get that $\mathbb{P}_{N}\big(y\in\mathcal{G}_{k}\big)\geq\phi_{L}e^{-4}p(N,L),$ where $\phi_{L}:=\mathbb{P}(\zeta_{y}\leq L-1)$, and we recall that $\zeta_{y}$ is the number of offspring of vertex $y$ and by (3.5) we have $D_{y}=\zeta_{y}+1$. Since the number of offspring are sampled independently for every parent in a BGW tree, this implies that $|\mathcal{G}_{k}|$ stochastically dominates a binomial random variable $B\sim\text{Bin}(N,e^{-4}\phi_{L}p(N,L)))$. By the Chernoff bound for a binomial distribution (see e.g. Theorem 2.21 in [22]), we get for $0<c<1$ that
@@
$\mathbb{P}_{N}\big(|\mathcal{G}_{k}|\geq ce^{-4}\phi_{L}Np(N,L)\ \text{ for all }\ k\leq e^{ce^{-4}\phi_{L}Np(N,L)}\big)\geq 1-\sum_{k=0}^{S}\mathbb{P}_{N}\big(|\mathcal{G}_{k}|<ce^{-4}\phi_{L}Np(N,L)\big).$ (7.4)
@@
Note that $(1-c)^{2}-2c>2c$ if $c<3-2\sqrt{2}$, and thus if we choose $c\in(0,3-2\sqrt{2})$ and set $c_{L}:=ce^{-4}\phi_{L}$ we have
$\sum_{k=0}^{S}\mathbb{P}_{N}(|\mathcal{G}_{k}|<c_{L}Np(N,L))\leq e^{-c_{L}Np(N,L)}.\qed$
@@ 7.2 Survival on a star
$\mathcal{G}_{k}^{\prime}=\mathcal{G}^{\prime}_{\rho,k}:=\{y\in\mathcal{G}_{k}:\ y\in\mathbf{C}^{*}_{kT}\},$ (7.5)
In words this is the set of good neighbours in the interval $J_{k}$, which are infected at time $kT$, and thus are infected at any time during $[kT,(k+2)T]$ since good neighbours do not recover during $J_{k}$. Furthermore, for any $0<s<t$, we denote by $\mathcal{F}_{s,t}$, the $\sigma$-algebra generated by the graphical construction up to time $s$, and the processes $\Delta^{\text{rec}}_{y}$ and $\Delta^{\text{up}}_{\{\rho,y\}}$ for all $y\in\mathcal{N}_{\rho}$ up to time $t$.
$\mathcal{W}_{k}:=\left\{|\mathcal{G}^{\prime}_{k}|\geq\delta\lambda Np(N,L)T\quad\text{and}\quad\int_{kT}^{(k+1)T}\mathds{1}_{\{\rho\notin\mathbf{C}_{s}^{*}\}}\mathrm{d}s\leq\frac{T}{2}\right\}.$
@@
(Local survival) Let $\lambda>0$ and $N\in\mathbb{N}$ be chosen such that $\tfrac{3}{2}\lambda T_{N}<1$ and set $\bar{k}=\lfloor e^{\delta\lambda^{2}T^{2}Np(N,L)/4}\rfloor$. Then there exists a large universal constant $C$ such that if $\lambda^{2}T^{2}N>C$, then, for any $k\leq e^{c_{L}Np(N,L)/2}$, we have
@@
The result follows analogously to [11, Proposition 3.10]. Note the condition $\tfrac{3}{2}\lambda T_{N}<1$, which does not appear in the statement of [11]. However, while reproducing the proof, we found that the argument only holds for sufficiently small $\lambda$, as noted in the proof of [11, Proposition 3.10]. In particular, in our setting, we require the condition $\tfrac{3}{2}\lambda T_{N}<1$. ∎
With this result we have control over the event that an infection persists for an exponential time if we already have a sufficiently large pool of infected good neighbours which can sustain the centre to be infected most of the time. It remains to estimate the probability to reach this amount of good infected neighbours in case that only the centre is initially infected.
@@
Let $\lambda>0$ and $N\in\mathbb{N}$ be chosen such that $2\lambda T_{N}<1$. Furthermore, let $s\leq e^{c_{L}Np(N,L)/2}$ and $k=\lfloor s/T\rfloor$. Then it holds that
@@
By the choice of $k$ we see that $kT\leq s\leq(k+1)T$, and thus the probability that $\rho$ does not recover in $[s,(k+2)T]$ is bounded from below by $e^{-2T}$. On the other hand, the probability that an infection event happens between a good neighbour $y\in\mathcal{G}_{k+2}$ and $\rho$ in the time interval $[s,(k+2)T]$ is bound from below by $1-e^{-\lambda T}\geq\frac{\lambda T}{2}$. Note that the last inequality holds since $\lambda T<1$. Now given that $\rho$ does not recover in $[s,(k+2)T]$ we conclude that on the event $\mathcal{S}_{\rho,N}\cap\{\rho\in\mathbf{C}^{*}_{s}\}$ the random variable $|\mathcal{G}^{\prime}_{k+2}|$ stochastically dominates a binomial random variable $B\sim\text{Bin}(\lceil 8\delta Np(N,L)\rceil,\tfrac{\lambda T}{2})$. Recall that $\delta=c_{L}/8$, where $c_{L}$ is the constant from Lemma 7.1. Therefore, we get that
@@
Next we can again use a Chernoff bound for the binomial distribution (see e.g. [22, Theorem 2.21]) to obtain
$\mathbb{P}_{\mathcal{T},N}\Big(B>\delta\lambda TNp(N,L)\Big)\geq 1-e^{-\delta\lambda TNp(N,L)}.$
This completes the proof. ∎